what exacty is residual income? How do we calculate it and what does it tell us?
Answer #1

Residual income = Net operating income - Required return on invested assets

.

Residual income is the excess of operating income over the expected return on the assets invested. Higher the residual income, better the performance. To understand this, let's take an example.

Your company takes up a project in which $100,000 worth of assets are invested. It is expected to make a return of 10%. However the net operating income from the project is $15,000. So, the residual income will be $5,000.

Residual income = Net Operating income - Required return

= 15,000 - (100,000*10%) = 15,000 - 10,000

= $5,000